1. Find the steady state temperature of the black body disc having radius 2cm and mass 7g at 450C surrounding temperature, when kept initially at 280C. Hence determine Stefan's constant.
1
Expert's answer
2021-08-10T17:55:16-0400

Solution:-

The disc gains heat "Q_{in}" due to thermal conductivity from the surrounding and losses heat "Q_{out}" due to radiation .

Thus, according to the Stefan Boltzmann law

Losses rate

"\frac{dQ_{out}}{dt}=\pi r^2\sigma(T_{sur}^3-T^4)\\\sigma=5.56\times10^{-8}W/m^2K^4\\"


"\frac{dQ_{out}}{dt}=\frac{d}{dt}(mc(T-T_{in}))=mc\frac{dT}{dt}"

Steady state

"\frac{dQ_{out}}{dt}=\frac{dQ_{in}}{dt}"

"\frac{dT}{dt}=\frac{\pi r^2\sigma T_{sur}^4}{mc}" "=\frac{\pi r^2\sigma T^4}{mc}"

Solution thi equation

"T(t)=\frac{1}{3\sqrt{{\frac{\pi r^2\sigma t}{mc}}+\frac{1}{(T_{start}-T_{sur})^3}}}+T_{sur}"

t=infinite

Put value

"T(t)=T_{sur}"


"T_{steady}=T_{sur}=273+45=318K=45\u00b0C"